Instructions and Grading Question 1. (3 points total) Below is a histogram of the salaries of the New York Yankees baseball team in 2011. Use the histogram to answer the following questions. Provided are some summary statistics of the data set. Standard Median Q1 Deviation Minimum Max Mean 03 $414,000 $32,000,000 $6,756,301 $8,468,058 $2,100,000 $559,625 $10,000,000 (1 point) Describe the histogram, commenting on the shape, center and overall spread. (Do your best at estimating what the center would be) a) Salaries of New York Yankees 2011 0 o 100 10 200 2 300 30 USD (n $100,000) (2 points) What proportion of players make more than $25,000,000? Show work. (note that 25,000,000 250). There is a total of 30 players. b) 100,000
Question 2. (7 points total) Based on the summary statistics in Question 1, answer the following. a.) (0.5 points) What is the Interquartile Range (IQR)? b.) (1.5 points) What are the cutoff values/fences for outliers? (Hint: use the formula (Q1-1.5*IQR) for the lower fence and (Q3+1.5*1QR) for the upper fence. c) (1 point) Based on your answer to part b, are there any outliers in the data? Why or why not? d.) (4 points) We see from the table that the maximum value in the data set is $32,000,000. Suppose that the next largest value is $22,000,000. Using this information and the table of summary statistics, draw a boxplot, making sure to label the minimum, Q1, median, Q3, maximum, and any outliers!
Answer #1

1. a) The shape of the histogram seems to be right-skewed. The mean can be computed as:

30 where Xi denotes the mid-value of ith bin and fi the frequency of the bin.

Thus, we get $7,000,000.

b). there is only one player above the 25,000,000 slab thus the proportion is 1/30=0.033

2. a). Here IQR is given as (Q3 - Q1)= $9,440,375

b). lower fence= Q1 - 1.5*IQR = -13,746,563

upper fence= Q3 + 1.5*IQR = $24,160,563

c). Yes, there is one outlier which isthe maximum value $32,000,000